on a winter day the temp. of the tap water is 20oc whereas the room temp. is 5oc. water is stored in a tank of capacity 0.5m^3 for household use. if it were possible to use the heat librated by the water to lift a 10kg mass vertically, how high can it be lifted as the water comes to the room temp. ? take g=10m/s^2

Ans=315km


Grade:10

1 Answers

nimo jar
6 years ago
change in temp of tank = \Delta t = 20 – 5 = 15                 [water will attain atmospheric temperature]
Heat liberated by water = Q = mc\Delta t = \deltaVc\Delta t          [c = specific heat capacity]
Q = 1 kg/m3 * 0.5 m3 * 4200 J/kg/oC * 15 oC = 31500 J
since, Q = mgh
thus, 31500 = 10 * 10 * h
h = 315 m